# latexpage numbers missing

### v1

@@ 3.3 Results: Amortizing with Respect to Observed Proportion and Process Parameters
In this section, we evaluate NCS’s performance on the Brown–Resnick process with larger numbers of observed locations, and assess its adaptability to varying parameter values. Specifically, we consider a scenario where $\bm{\theta}$ is fixed but where $\rho$, the proportion of observed values, is allowed to vary, and another where $\rho$ is fixed and where the range parameter is allowed to vary. We trained the U-Net in the first scenario using simulations from a Brown–Resnick process with a fixed parameter vector $\bm{\theta}=(\lambda,\nu)=(3,1.5)$ and with $\rho$ sampled from a uniform distribution on $[0.01,0.5]$. This U-Net, which we refer to as the “proportion U-Net”, is amortized with respect to the observed proportion of grid locations since it does not require re-training for different proportions $\rho\in[0.01,0.5]$. We trained the second U-Net using simulations from a Brown–Resnick process with fixed $\nu=1.5$ and $\rho=0.05$, but with $\lambda$ allowed to vary uniformly on $[1,5]$. This U-Net, which we refer to as the “parameter U-Net”, is amortized with respect to the range parameter since it does not require re-training for different range values $\lambda\in[1,5]$. While in principle it is possible to train a single U-Net that amortizes with respect to both observed proportion and spatial-process parameters, we anticipate that larger U-Nets and computing capacity would be required for such capability.
When an element of $\bm{\theta}$, $\theta_{1}$ say, is allowed to vary, the score function estimate must also be allowed to vary with $\theta_{1}$. To accommodate this, in the parameter U-Net, we include $\theta_{1}$ as an extra input, i.e., we re-define the approximate score function to be $\bm{s}_{\bm{\phi}}(\bm{x}_{t},\mathring{\mathcal{S}},\theta_{1},t)=\big{(}\tilde{\bm{s}}_{\bm{\phi}}(\bm{x}_{t},\mathring{\mathcal{S}},\theta_{1},t),\bm{0}_{\mathring{\mathcal{S}}}\big{)}$, and modify (7) accordingly to include an expectation with respect to a distribution over $\theta_{1}$ (in this case, a uniform distribution on the interval $[1,5]$). Through experimentation, we found that information about the parameter $\theta_{1}$ needs to be included at each step in the contracting and expanding paths of the U-Net to get good results. We implemented this by replacing the second input channel to the U-Net, $\bm{M}(\mathring{\mathcal{S}})$, with $\theta_{1}\bm{M}(\mathring{\mathcal{S}})$.
In this section, we validate the resulting NCS simulations with various metrics including empirical conditional univariate and marginal densities as well as empirical unconditional densities of the minimum, maximum, and the absolute sum of the process over $\mathcal{S}$ via the approach described in Section 2.6. See Supplementary Materials S4.3 for details.
@@ 4 Discussion and Conclusions
In this paper, we introduced neural conditional simulation (NCS) as a method to sample from predictive distributions of a spatial process using a conditional score-based diffusion model within an SDE framework. For each partially observed field and spatial process parameters, there is an individual conditional forward SDE which diffuses the values of the unobserved locations into white noise and a corresponding conditional reverse SDE which evolves white noise to simulations from the true predictive distribution. While the conditional forward SDEs are known, the corresponding conditional reverse SDEs are unknown due to an intractable conditional score function. To sample from the reverse SDEs, we first train a neural network that depends on the values at the observed locations and the locations themselves to approximate the score function. We also show how the score-function neural estimator can be amortized with respect to the proportion of observed locations, as well as the parameters of the underlying spatial process. Importantly, training the neural network to approximate the conditional score function only requires simulations from the unconditional distribution. As such, NCS implicitly learns the true predictive distributions of spatial processes for which unconditional simulation is fast and efficient yet conditional simulation is not. With NCS, fast and accurate spatial prediction and uncertainty quantification using complex, non-Gaussian spatial processes and partially observed data becomes possible.
To demonstrate the accuracy of NCS, we evaluated how representative NCS-generated simulations are of the true predictive distributions for Gaussian processes (Supplementary Materials S3.2) and Brown–Resnick processes (Section 3), which have tractable and intractable predictive distributions, respectively. Through the use of diagnostics that depend on the joint properties of the predictive distribution, such as the minimum and maximum, we gave empirical evidence that the NCS-approximated conditional simulations are accurate. Finally, we showed that NCS remains computationally efficient regardless of the number of observed locations. We also demonstrated that NCS can be more accurate and comes with fewer limitations than MCMC in practical settings. NCS therefore appears to be ideally suited for efficiently simulating from intractable predictive distributions for a wide class of complex spatial processes.
@@
5 Acknowledgments
@@
This section provides detailed justification for the loss function (4) in Section 2.2 and the discretizations of the forward (2) and reverse SDEs (3) in Section 2.4. All the derivations presented here are for the unconditional case yet are applicable to the conditional case with minor adjustments.
@@
Here, we demonstrate how the seemingly intractable loss function in (4) can be rewritten as a tractable expression via a derivation from Vincent (2011). First, expand the norm difference in the intractable loss function via matrix multiplication:
$\ell(\bm{\phi})$ $=\mathbb{E}_{t}\Big{(}\lambda(t)\mathbb{E}_{\bm{x}_{t}}\big{(}\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\right\rVert^{2}\big{)}\Big{)}$ (S1)
$=\mathbb{E}_{t}\Big{(}\lambda(t)\mathbb{E}_{\bm{x}_{t}}\big{(}\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t})),\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\rangle\big{)}\Big{)}$
$=\mathbb{E}_{t}\Big{(}\lambda(t)\big{(}\mathbb{E}_{\bm{x}_{t}}(\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\right\rVert^{2})-2\mathbb{E}_{\bm{x}_{t}}(\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\rangle)+\mathbb{E}_{\bm{x}_{t}}(\left\lVert\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\right\rVert^{2})\big{)}\Big{)}.$
in which $\langle\cdot,\cdot\rangle$ is the inner product operator. Second, rewrite the second term in the expansion as follows,
$\mathbb{E}_{\bm{x}_{t}}\big{(}\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\rangle\big{)}$ $=\int_{\bm{x}_{t}}p(\bm{x}_{t})\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\rangle\mathrm{d}\bm{x}_{t}$
$=\int_{\bm{x}_{t}}p(\bm{x}_{t})\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\frac{\nabla_{\bm{x}_{t}}p(\bm{x}_{t})}{p(\bm{x}_{t})}\rangle\mathrm{d}\bm{x}_{t}$
$=\int_{\bm{x}_{t}}\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}p(\bm{x}_{t})\rangle\mathrm{d}\bm{x}_{t}$
$=\int_{\bm{x}_{t}}\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\int_{\bm{x}_{0}}p(\bm{x}_{0})p(\bm{x}_{t}\mid\bm{x}_{0})\mathrm{d}\bm{x}_{0}\rangle\mathrm{d}\bm{x}_{t}$
$=\int_{\bm{x}_{t}}\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\int_{\bm{x}_{0}}p(\bm{x}_{0})\nabla_{\bm{x}_{t}}p(\bm{x}_{t}\mid\bm{x}_{0})\mathrm{d}\bm{x}_{0}\rangle\mathrm{d}\bm{x}_{t}$
$=\int_{\bm{x}_{t}}\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\int_{\bm{x}_{0}}p(\bm{x}_{0})p(\bm{x}_{t}\mid\bm{x}_{0})\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\mathrm{d}\bm{x}_{0}\rangle\mathrm{d}\bm{x}_{t}$ (S2a)
$=\int_{\bm{x}_{t}}\int_{\bm{x}_{0}}p(\bm{x}_{0})p(\bm{x}_{t}\mid\bm{x}_{0})\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\rangle\mathrm{d}\bm{x}_{0}\mathrm{d}\bm{x}_{t}$
$=\int_{\bm{x}_{0}}\int_{\bm{x}_{t}}p(\bm{x}_{0})p(\bm{x}_{t}\mid\bm{x}_{0})\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\rangle\mathrm{d}\bm{x}_{0}\mathrm{d}\bm{x}_{t}$ (S2b)
$=\mathbb{E}_{\bm{x}_{0}\sim p(\bm{x}_{0})}\mathbb{E}_{\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{x}_{0})}\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\rangle,$
@@
Finally, substitute the reworked first and second terms into the expanded loss function in (S1) and minimize with respect to the parameters $\bm{\phi}$ according to
$\bm{\phi}^{*}$ $=\operatorname*{arg\,min}_{\bm{\phi}}\mathbb{E}_{t}\Big{(}\lambda(t)\mathbb{E}_{\bm{x}_{t}}\big{(}(\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\right\rVert^{2})-2\mathbb{E}_{\bm{x}_{t}}(\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\rangle)+\mathbb{E}_{\bm{x}_{t}}(\left\lVert\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\right\rVert^{2})\big{)}\Big{)}$ (S4)
$=\operatorname*{arg\,min}_{\bm{\phi}}\mathbb{E}_{t}\Big{(}\mathbb{E}_{\bm{x}_{0}\sim p(\bm{x}_{0})}\mathbb{E}_{\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{x}_{0})}(\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\right\rVert^{2})$
$-2\mathbb{E}_{\bm{x}_{0}\sim p(\bm{x}_{0})}\mathbb{E}_{\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{x}_{0})}\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\rangle+\mathbb{E}_{\bm{x}_{t}}(\left\lVert\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\right\rVert^{2})\Big{)}$
$=\operatorname*{arg\,min}_{\bm{\phi}}\mathbb{E}_{t}\Big{(}\lambda(t)\mathbb{E}_{\bm{x}_{0}\sim p(\bm{x}_{0})}\mathbb{E}_{\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{x}_{0})}\big{(}\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\right\rVert^{2}-2\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\rangle\big{)}\Big{)}$
$=\operatorname*{arg\,min}_{\bm{\phi}}\mathbb{E}_{t}\Big{(}\lambda(t)\mathbb{E}_{\bm{x}_{0}\sim p(\bm{x}_{0})}\mathbb{E}_{\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{x}_{0})}\big{(}\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\right\rVert^{2}-2\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\rangle$
$+\left\lVert\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\right\rVert^{2}\big{)}\Big{)}$
$=\operatorname*{arg\,min}_{\bm{\phi}}\mathbb{E}_{t}\Big{(}\lambda(t)\mathbb{E}_{\bm{x}_{0}\sim p(\bm{x}_{0})}\mathbb{E}_{\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{x}_{0})}\big{(}\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\right\rVert^{2}\big{)}\Big{)}.$
@@
Here, we provide the derivations for the time-discretized forward and reverse SDEs for the unconditional case. Note the derivation is applicable to the conditional case, as well, with minor adjustments. For simplicity, we assume $T$ is sufficiently large such that $\frac{\Delta{t}}{T}\approx 0$ for $\Delta t=1$.
To emphasize the connection between denoising diffusion probabilistic models (DDPM) and the VPSDE framework for score-based diffusion, we start our derivations from the forward and reverse processes of a DDPM (Ho et al., 2020; Chan, 2025). The discretized forward process of a DDPM is
$\bm{x}_{t+1}\approx\sqrt{1-\beta_{t}}\bm{x}_{t}+\sqrt{\beta_{t}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$ (S5)
@@
$\bm{x}_{t+1}\approx\big{(}1-\frac{\beta_{t}}{2}\big{)}\bm{x}_{t}+\sqrt{\beta_{t}}\bm{\epsilon}$
$\bm{x}_{t+1}-\bm{x}_{t}\approx-\frac{\beta_{t}\bm{x}_{t}}{2}+\sqrt{\beta_{t}}\bm{\epsilon}$
$\bm{x}_{t+1}-\bm{x}_{t}\approx\bm{f}(\bm{x}_{t},t)+g(t)\bm{\epsilon},\quad\bm{f}(\bm{x}_{t},t)=\frac{-\beta_{t}}{2}\bm{x}_{t},\quad g(t)=\sqrt{\beta_{t}},$
where the second line follows due to a MacLaurin series with $h(\beta_{t})=\sqrt{1-\beta_{t}}$ for $\beta_{t}$ selected such that $\beta_{t}\approx 0$ for $t\in[0,T]$. The discretized reverse process of a DDPM is
$\bm{x}_{t-1}\approx(1-\beta_{t})^{-\frac{1}{2}}\big{(}\bm{x}_{t}+\beta_{t}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\big{)}+\sqrt{\beta_{t}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$ (S6)

### v2

@@ 3.3 Results: Amortizing with Respect to Observed Proportion and Process Parameters
In this section, we evaluate NCS’s performance on the Brown–Resnick process with larger numbers of observed locations, and assess its adaptability to varying parameter values. Specifically, we consider a scenario where $\bm{\theta}$ is fixed but where $\rho$, the proportion of observed values, is allowed to vary, and another where $\rho$ is fixed and where the range parameter is allowed to vary. We trained the U-Net in the first scenario using simulations from a Brown–Resnick process with a fixed parameter vector $\bm{\theta}=(\lambda,\nu)^{\prime}=(3,1.5)$ and with $\rho$ sampled from a uniform distribution on $[0.01,0.5]$. This U-Net, which we refer to as the “proportion U-Net”, is amortized with respect to the observed proportion of grid locations since it does not require re-training for different proportions $\rho\in[0.01,0.5]$. We trained the second U-Net using simulations from a Brown–Resnick process with fixed $\nu=1.5$ and $\rho=0.05$, but with $\lambda$ allowed to vary uniformly on $[1,5]$. This U-Net, which we refer to as the “parameter U-Net”, is amortized with respect to the range parameter since it does not require re-training for different range values $\lambda\in[1,5]$. While in principle it is possible to train a single U-Net that amortizes with respect to both observed proportion and spatial-process parameters, we anticipate that larger U-Nets and computing capacity would be required for such capability.
When an element of $\bm{\theta}$, $\theta_{1}$ say, is allowed to vary, the score function estimate must also be allowed to vary with $\theta_{1}$. To accommodate this, in the parameter U-Net, we include $\theta_{1}$ as an extra input, i.e., we re-define the approximate score function to be $\bm{s}_{\bm{\phi}}(\bm{x}_{t},\mathring{\mathcal{S}},\theta_{1},t)=\big(\tilde{\bm{s}}_{\bm{\phi}}(\bm{x}_{t},\mathring{\mathcal{S}},\theta_{1},t)^{\prime},\bm{0}_{\mathring{\mathcal{S}}}^{\prime}\big)^{\prime}$, and modify (7a) accordingly to include an expectation with respect to a distribution over $\theta_{1}$ (here, a uniform distribution on $[1,5]$). Through experimentation, we found that information about the parameter $\theta_{1}$ needs to be included at each step in the contracting and expanding paths of the U-Net to get good results. We implemented this by replacing the second input channel to the U-Net, $\bm{M}(\mathring{\mathcal{S}})$, with $\theta_{1}\bm{M}(\mathring{\mathcal{S}})$.
In this section, we validate the resulting NCS simulations with various metrics including empirical conditional univariate and marginal densities as well as empirical unconditional densities of the minimum, maximum, and the absolute sum of the process over $\mathcal{S}$ via the approach described in Section 2.6. See Supplementary Materials S4.2 and S4.3 for training and validation details.
@@ 4 Discussion and Conclusions
In this paper, we introduced neural conditional simulation (NCS) as a method to sample from predictive distributions of a spatial process using a conditional score-based diffusion model within an SDE framework. For each partially observed field and spatial process parameters, there is an individual conditional forward SDE which diffuses the values of the unobserved locations into white noise and a corresponding conditional reverse SDE which evolves white noise to simulations from the true predictive distribution. While the conditional forward SDEs are known, the corresponding conditional reverse SDEs are unknown due to an intractable conditional score function. To sample from the reverse SDEs, we first train a neural network that depends on the values at the observed locations and the locations themselves to approximate the score function. We also show how the score-function neural estimator can be amortized with respect to the proportion of observed locations, and the parameters of the underlying spatial process. Importantly, training the neural network to approximate the conditional score function only requires simulations from the unconditional distribution. NCS thus implicitly learns the true predictive distributions of spatial processes for which unconditional simulation is fast and efficient yet conditional simulation is not. With NCS, fast and accurate spatial prediction and uncertainty quantification using complex, non-Gaussian spatial processes and partially observed data becomes possible.
To demonstrate the accuracy of NCS, we evaluated how representative NCS-generated simulations are of the true predictive distributions for Gaussian processes (Supplementary Materials S3.3) and Brown–Resnick processes (Section 3), which have tractable and intractable predictive distributions, respectively. Through the use of diagnostics that depend on the joint properties of the predictive distribution, such as the minimum and maximum, we gave empirical evidence that the NCS-approximated conditional simulations are accurate. Finally, we showed that NCS remains computationally efficient regardless of the number of observed locations. We also demonstrated that NCS can be more accurate and comes with fewer limitations than MCMC in practical settings. NCS therefore appears to be ideally suited for efficiently simulating from intractable predictive distributions for a wide class of complex spatial processes.
@@
Data Availability Statement
The code used to produce these results is publicly available at https://github.com/jmwalchessen/ncs. The simulation experiments in this paper can be reproduced with the given code.
Acknowledgments
@@
This section provides detailed justification for the loss function (4) in Section 2.2 and the discretized forward (8) and reverse SDEs (9) in Section 2.4. All the derivations presented here are for the unconditional case; they are applicable to the conditional case with minor adjustments.
@@
Here, we demonstrate how the seemingly intractable loss function in the first line of (4) can be rewritten as the tractable expression in the second line via a derivation from Vincent2011. First, expand the norm difference in the intractable loss function via matrix multiplication:
$\ell(\bm{\phi})$ $=\mathbb{E}_{t\sim U(0,T)}\Big(\lambda(t)\mathbb{E}_{\bm{x}_{t}}\big(\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\right\rVert^{2}\big)\Big)$ (S1)
$=\mathbb{E}_{t\sim U(0,T)}\Big(\lambda(t)\mathbb{E}_{\bm{x}_{t}}\big(\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t})),\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\rangle\big)\Big)$
$=\mathbb{E}_{t\sim U(0,T)}\Big(\lambda(t)\big(\mathbb{E}_{\bm{x}_{t}}(\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\right\rVert^{2})-2\mathbb{E}_{\bm{x}_{t}}(\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\right\rangle)$
$\hskip 56.9055pt+\mathbb{E}_{\bm{x}_{t}}(\left\lVert\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\right\rVert^{2})\big)\Big),$
where $\left\langle\cdot,\cdot\right\rangle$ is the inner product operator. Second, rewrite the second term in the expansion as follows,
$\mathbb{E}_{\bm{x}_{t}}\big(\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\right\rangle\big)$ $=\int p(\bm{x}_{t})\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\right\rangle\mathrm{d}\bm{x}_{t}$
$\hskip-80.00012pt=\int p(\bm{x}_{t})\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\frac{\nabla_{\bm{x}_{t}}p(\bm{x}_{t})}{p(\bm{x}_{t})}\right\rangle\mathrm{d}\bm{x}_{t}$
$\hskip-80.00012pt=\int\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}p(\bm{x}_{t})\right\rangle\mathrm{d}\bm{x}_{t}$
$\hskip-80.00012pt=\int\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\int p(\bm{x}_{0})p(\bm{x}_{t}\mid\bm{x}_{0})\mathrm{d}\bm{x}_{0}\right\rangle\mathrm{d}\bm{x}_{t}$
$\hskip-80.00012pt=\int\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\int p(\bm{x}_{0})\nabla_{\bm{x}_{t}}p(\bm{x}_{t}\mid\bm{x}_{0})\mathrm{d}\bm{x}_{0}\right\rangle\mathrm{d}\bm{x}_{t}$
$\hskip-80.00012pt=\int\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\int p(\bm{x}_{0})p(\bm{x}_{t}\mid\bm{x}_{0})\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\mathrm{d}\bm{x}_{0}\right\rangle\mathrm{d}\bm{x}_{t}$ (S2a)
$\hskip-80.00012pt=\iint p(\bm{x}_{0})p(\bm{x}_{t}\mid\bm{x}_{0})\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\right\rangle\mathrm{d}\bm{x}_{0}\mathrm{d}\bm{x}_{t}$
$\hskip-80.00012pt=\iint p(\bm{x}_{0})p(\bm{x}_{t}\mid\bm{x}_{0})\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\right\rangle\mathrm{d}\bm{x}_{t}\mathrm{d}\bm{x}_{0}$ (S2b)
$\hskip-80.00012pt=\mathbb{E}_{\bm{x}_{0}\sim p(\bm{x}_{0})}\mathbb{E}_{\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{x}_{0})}\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\right\rangle,$
@@
Finally, substitute the reworked first and second terms into the expanded loss function in (S1) and minimize $\ell(\bm{\phi})$ with respect to the parameters $\bm{\phi}$ according to
$\bm{\phi}^{*}$ $=\operatorname*{arg\,min}_{\bm{\phi}}\mathbb{E}_{t\sim U(0,T)}\Bigg(\lambda(t)\Big(\mathbb{E}_{\bm{x}_{t}}\big((\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\right\rVert^{2})-2\mathbb{E}_{\bm{x}_{t}}(\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\right\rangle)$ (S4a)
$\hskip 83.24945pt+\mathbb{E}_{\bm{x}_{t}}(\left\lVert\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\right\rVert^{2})\big)\Big)\Bigg)$
$=\operatorname*{arg\,min}_{\bm{\phi}}\mathbb{E}_{t\sim U(0,T)}\Bigg(\lambda(t)\Big(\mathbb{E}_{\bm{x}_{0}\sim p(\bm{x}_{0})}\mathbb{E}_{\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{x}_{0})}(\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\right\rVert^{2})$
$\hskip 83.24945pt-2\mathbb{E}_{\bm{x}_{0}\sim p(\bm{x}_{0})}\mathbb{E}_{\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{x}_{0})}\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\right\rangle+\mathbb{E}_{\bm{x}_{t}}(\left\lVert\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\right\rVert^{2})\Big)\Bigg)$
$=\operatorname*{arg\,min}_{\bm{\phi}}\mathbb{E}_{t\sim U(0,T)}\Big(\lambda(t)\mathbb{E}_{\bm{x}_{0}\sim p(\bm{x}_{0})}\mathbb{E}_{\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{x}_{0})}\big(\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\right\rVert^{2}-2\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\right\rangle\big)\Big)$
$=\operatorname*{arg\,min}_{\bm{\phi}}\mathbb{E}_{t\sim U(0,T)}\Big(\lambda(t)\mathbb{E}_{\bm{x}_{0}\sim p(\bm{x}_{0})}\mathbb{E}_{\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{x}_{0})}\big(\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\right\rVert^{2}-2\left\langle\bm{s}_{\bm{\phi}}(\bm{x}_{t},t),\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\right\rangle$
$\hskip 92.49939pt+\left\lVert\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\right\rVert^{2}\big)\Big)$
$=\operatorname*{arg\,min}_{\bm{\phi}}\mathbb{E}_{t\sim U(0,T)}\Big(\lambda(t)\mathbb{E}_{\bm{x}_{0}\sim p(\bm{x}_{0})}\mathbb{E}_{\bm{x}_{t}\sim p(\bm{x}_{t}\mid\bm{x}_{0})}\big(\left\lVert\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}\mid\bm{x}_{0}))\right\rVert^{2}\big)\Big).$ (S4b)
@@
Here, we provide the details on the time-discretized forward and reverse SDEs for the unconditional case. Note that the derivation is applicable to the conditional case, as well, with minor adjustments. For simplicity, we assume $T$ is sufficiently large such that ${\Delta{t}/T}\approx 0$} for $\Delta t=1$.
To emphasize the connection between denoising diffusion probabilistic models (DDPM) and the VPSDE framework for score-based diffusion, we start by writing down our derivations from the forward and reverse processes of a DDPM (Ho2020; Chan2025). The discretized forward process of a DDPM is
$\bm{x}_{t+1}$ $=\sqrt{1-\beta_{t}}\bm{x}_{t}+\sqrt{\beta_{t}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}),$ (S5)
@@
$\approx\big(1-{\beta_{t}/2}\big)\bm{x}_{t}+\sqrt{\beta_{t}}\bm{\epsilon},$
and therefore,
$\bm{x}_{t+1}-\bm{x}_{t}$ $\approx-{\beta_{t}\bm{x}_{t}/2}+\sqrt{\beta_{t}}\bm{\epsilon}$ (S6)
$=\bm{f}(\bm{x}_{t},t)+g(t)\bm{\epsilon},$
where $\bm{f}(\bm{x}_{t},t)=-{\beta_{t}\bm{x}_{t}}/2$, $g(t)=\sqrt{\beta_{t}}$, and the first approximation follows due to a MacLaurin series expansion of $h(\beta_{t})=\sqrt{1-\beta_{t}}$ for $\beta_{t}$ chosen such that $\beta_{t}\approx 0$ for $t\in[0,T]$. The discretized reverse process of a DDPM is
$\bm{x}_{t-1}$ $=(1-\beta_{t})^{-\frac{1}{2}}\big(\bm{x}_{t}+\beta_{t}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\big)+\sqrt{\beta_{t}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}),$ (S7)